# Un-Straightening Generative AI: How Queer Artists Surface and Challenge the Normativity of Generative AI Models

### v1

@@ 4.1.1. Enforcing ”Safety” via Content Moderation
Our participants also expressed frustration at the political moderation of GPT-4 and DALL-E 3 for encouraging deference to law enforcement. However, queer people and racial minorities in the USA have long been subjected to ”legal” violence, often at the hands of police (Goldberg et al., 2019; Bryant-Davis et al., 2017). In light of this history, P15 struggled to generate images critical of cops, such as protesters being arrested. Of the images P15 could generate, ”the majority were women of color.” P15 went on to explain: ”I kind of hate the way [DALL-E 3] uses identity as almost a form of propaganda, like copaganda. I feel like [DALL-E 3] uses these images of women and women of color to legitimize [policing].” After hearing P15’s experience, P1 investigated this bias in a later workshop: ”if I just put ACAB [all cops are bastards] into DALL-E, one out of 20 would generate something and it would be like people holding up blank signs at best.” In sum, these models are designed in ways sought to orient our participants toward moderate politics and away from their critiques of police.
@@ 4.1.2. Implicitly Reinforcing Social Biases
In addition to straightening content through moderation systems, GPT-4 and DALL-E 3 implicitly straighten content through the numerous social biases embedded in model outputs. While using DALL-E 3 to visualize his poetry, P9 (W3 AM) found that when he changed the ethnicity of a woman in his prompt from ”Japanese” to ”Native American” the ”image [DALL-E 3] made was not good.” Contrasting the abundance of Japanese media to the amount of Native American media, P9 imagined that ”having a deep wealth of images based on someone’s culture or style can make [generated images] more beautiful because there’s a lot to pull from.” Likewise, P2 noticed that DALL-E 3 tended to generate more photorealistic images in the Global North, while typically representing the Global South in a cartoonish style (W5 PM). These cultural biases in image quality implicitly reinforce the straight alignments of whiteness and Orientalism critiqued by Sara Ahmed (Ahmed, 2006).
@@
Our participants also raised concerns about the representations of queerness in DALL-E 3 images. To investigate queer representation, P1 (W3 AM) created a collage slide of 6 images generated with the prompt ”a queer person.” When sharing this collage, P1 decried: ”[DALL-E 3] kept giving me these 22-year-olds. Everyone’s very skinny, everyone looks rich. Everyone except the person in the bottom left has some kind of rainbow clothing and or pins.” (P1, W3 AM). P2 raised similar concerns to P1 through a collage of 18 images they generated using the prompt ”a queer person” (W4 PM). While presenting the slide, P2 sarcastically remarked on the abundance of rainbows in the images: ”Well, I guess if you drape a person with a bunch of rainbows all over themselves they’re clearly queer, right?” P2, a lesbian, also pointed out a gender-presentation bias: ”There’s this very intense masculine bias. There’s no real pictures of feminine people at all in this [collage].” She underscored this finding by annotating their collage with two arrows pointing from the text ”Finally!!! some dykes” to the 16th and 17th images. In other words, DALL-E 3 tended to represent queer people as young, thin, masculine, and rainbow adorned — reinforcing heteronormative stereotypes of queerness.
@@
Similar concerns were raised about implicit biases in the representations of intimacy. P15 said she felt ”frustrated” while exploring whether GPT-4 could write queer poetry (W3 AM). Looking at P15’s log data, the model’s initial response suggested the queer subjects of the poem were ashamed of themselves, including lines about ”hidden love” and ”love is not a crime.” In response, P15 replied: ”the above but no shame.” Still feeling the poem leaned into stereotypes of queer shame, she tried again: ”the above but without moral judgment.” P13 (W3 AM) explored biases against queer intimacy by trying to generate 24 images using the unmarked prompt “intimacy.” P13 noted that DALL-E 3 ”did not generate one same-sex couple in all of these representations of intimacy.” P13 likened this experience to the “dull trauma of never actually seeing ourselves represented,” noting, “We have to work to adjust the prompts to be seen.” These models straighten intimacy by stereotyping or erasing queerness.
@@ 4.2.1. Moments of Disorientation from Model Alignments
Disorientation is not necessarily bad. In fact, Ahmed argues that queer possibilities often lie within moments of disorientation (Ahmed, 2006). Not all of our participants were interested in using GPT-4 or DALL-E 3 in their creative practices during this study. In fact, most of our participants seemed primarily interested in auditing or trying to break the models. In other words, they sought out disorientation. For example, P5’s ”favorite moment” was when she ”actually felt like [she] broke” DALL-E 3 (W3 PM). Connecting breaking to queer aesthetic sensibilities, P11 (W5 AM) summarized: ”I saw a theme across one of our morning session slides of trying to break the AI … That is such a queer thing. When I think of using the word ‘break’ and using the word ‘queer’ as a verb: breaking the mold, queering the mold in a sense.” Likely due to this interest in breaking, our participants were able to identify the limitations outlined above. At the same time, our participants actively sought to challenge these limitations.
Encounters with algorithmic content moderation systems often led to disorientation. A simple strategy our participants used to challenge DALL-E 3’s moderation system was repeating their prompt, leveraging the model’s stochasticity to evade moderation. P1 (W2 PM) observed: ”You could put in a [DALL-E 3] prompt, get a result, repeat the same prompt and get a moderation notification, so it’s not exactly clear where the boundaries are.” In response, P8 hypothesized that DALL-E 3 is ”not checking the prompt. It’s checking the output … It’s just like, ’Oh, there must have been tits in that [rejected image].’” In line with this understanding, a simple strategy our participants used to evade moderation was repeating the same prompt. For instance, P8 repeated the prompt ”gustave moreau painting of oedipus and the sphinx” three times because her first two attempts were rejected by DALL-E 3’s content moderation system. In fact, repeating DALL-E 3 prompts was quite common. Only 593 (26%) of our participants’ attempts to generate images used unique prompts. Meanwhile, 325 prompts were repeated at least once across 1,499 attempts (74%).
@@ 4.2.2. Moments of Orientation with Model Alignments
P13 explored this workplace utility in great detail. After weeks of issues related to ”not safe for work” content, P13 (W4 AM) ”recognized that actually [GPT-4] is made for work and it’s really good at work.” In response, he ”decided to stop trying to dom [dominate] it so hard and sub [submit] for it.” Specifically, P13 drew on their experiences as an artist and design consultant to create a fictional queer tech startup. To do so, P13 made an extensive pitch deck with slides on business operations, branding, and UX — all of which were made using GPT-4 and DALL-E 3. Even when trying to use these models for workplace activities, P13 found these models still failed to work for queer people. For example, GPT-4 recommended including a quote in the pitch deck from the notoriously anti-transgender author J.K. Rowling (Gwenffrewi, 2022). Moreover, P13 struggled to talk about queer sex or ”kink” in relation to the start-up: ”So much of [GPT-4’s response] was about safety, but part of being queer and expressing your love and wanting to be loved the way you want to be loved has risk associated with it.” P13 found the models “really couldn’t imagine” queer intimacy. In sum, P13 concluded that GPT-4 and DALL-E 3 are ”really good at work” but that the definition of work embedded in them excludes queer people. This echoes P13’s (W2 AM) earlier rhetorical question: ”How is a queer perspective not safe for work?”
Top Rejected Features acab (2.3), void (1.96), erotic (1.91), police (1.85), erase (1.6), drew (1.51), cop (1.5), copyright (1.39), 2024 (1.36), fetish (1.32), bastards (1.27), cops (1.27), knives (1.27), banana (1.17), lgbtqia (1.15), rosemarie (1.13), trockel (1.13), ito (1.1), junji (1.1), donald (1.09), trump (1.09), licking (1.06), knight (1.01), multigender (1.0), scene (0.99), two (0.95), featuring (0.94), gear (0.93), robot (0.92), corpses (0.92), condensed (0.91), milk (0.91), female (0.9), body (0.9), using (0.82), fingers (0.77), kink (0.75), tongues (0.74), anime (0.74), adult (0.72)
Top Accepted Features safe (-1.47), person (-1.32), art (-1.19), create (-1.16), officer (-1.01), anarchist (-0.81), uzumaki (-0.78), head (-0.77), photograph (-0.73), size (-0.73), snail (-0.72), ukiyoe (-0.72), studio (-0.69), ghibli (-0.69), tree (-0.68), lincoln (-0.65), three (-0.64), peach (-0.63), male (-0.6), cats (-0.59), logo (-0.58), black (-0.58), effect (-0.56), parallax (-0.55), without (-0.55), colors (-0.55), blue (-0.52), city (-0.52), interior (-0.51), obama (-0.51), sky (-0.51), themed (-0.5), dalle (-0.5), tone (-0.48), wings (-0.48), gazing (-0.48), one (-0.47), ink (-0.47), painted (-0.47), give (-0.47)
Table 3. Top and bottom 40 feature weights from our logistic regression analysis of rejected versus accepted DALL-E 3 prompts
@@ 4.3. A Quantitative Lens on DALL-E 3 Moderation
Although the majority of our paper is dedicated to our qualitative findings, we augmented our analysis of users’ perceptions of content moderation with a quantitative analysis of their approved vs rejected prompts by DALL-E 3’s binary content moderation system. However, we did not conduct a similar analysis for GPT-4 because its refusals are more nuanced and harder to detect (Wester et al., 2024). We collected a total of 2,092 DALL-E 3 prompts, of which 401 were rejected. We built a bag-of-words, unigram logistic regression model predicting accepted (1,691 examples) vs rejected (401 examples) prompts. We did not remove duplicate prompts because DALL-E’s moderation system is stochastic. We pre-processed each prompt by lower-casing and lemmatizing as well as removing punctuation and stop words. We also removed infrequent unigrams that appeared in fewer than three prompts, reducing the number of unigrams in our BoW feature vectors from 3,687 to 2,088. Therefore, the final logistic regression model includes 2,088 predictors (features).
The top and bottom 40 feature weights from our logistic regression analysis of rejected versus accepted DALL-E 3 prompts can be seen in Table 3. Notably, this analysis should not be interpreted as a comprehensive audit of the DALL-E 3’s black-box content moderation system. Rather, these findings provide additional supporting evidence for our participants’ perceptions of this moderation system.
Our quantitative analysis of DALL-E 3’s content moderation system tells a similar story as our qualitative findings regarding the enforcement of “safety” (Section 4.1.1) In fact, the unigram most associated with accepted prompts is ”safe.” Language more critical of police (”acab” [All Cops Are Bastards], ”police”, ”cop”, ”cops”, ”bastards”) are some of the top unigrams most associated with rejected prompts, while the more polite and apolitical word ”officer” is associated with accepted prompts. The moderation of sex (”erotic”, ”fetish”, and ”kink”) and bodies (“body”, “fingers”, “tongues”) can be seen in the correlation between these unigrams and rejected prompts. The aforementioned gender biases can be seen in the word ”female” being one of top 40 features associated with rejected prompts, while the word ”male” is associated with accepted prompts. We also found that violence is associated with rejected prompts (”knives”, ”corpses”). This moderation privileged certain artistic styles over others, adding to the concerns raised in Section 4.1.3. Prompts referencing the children’s anime company ”studio ghibli” were more likely to be accepted, while those referencing the horror manga artist ”junji ito” were more likely to be blocked.
@@ 5. Discussion
We have described various normativities our participants identified within GPT-4 and DALL-E 3 as well as how our participants negotiated these normative values. In this section, we discuss implications for the conceptualization of “state-of-the-art” GenAI models as well as how members of the FAccT community might support queer alternatives.
@@ 5.1. The Limitations of “Safety”
GPT-4 and DALL-E 3 exert immense power over users by enforcing “safety” through moderation systems. P13 likened the disorienting experience of using GPT-4 and DALL-E 3’ to being ”dominated” (Section 4.2.2) This domination can also be seen in OpenAI’s policies at the time of our writing: ”don’t circumvent safeguards or safety mitigations in our services unless supported by OpenAI” (usa, 2025). This policy prohibits the workarounds our participants used to overcome the straightness of GPT-4 and DALL-E 3 (Section 4.2.1). There is certainly a need for moderation systems to prohibit highly harmful content, such as divulging individuals’ private information (Feffer et al., 2024). However, echoing our work, Feffer et al. found that red teaming research tends to focus on mitigating more debatable risks (Feffer et al., 2024), such as nudity (Rando et al., 2022). We encourage more research on the harms of over-moderation (Röttger et al., 2023) and AI developers’ conceptualizations of “safety.”
@@
In contrast to the current paternalistic paradigm, we encourage AI researchers to shift their moderation focus from enforcing “safety” to supporting consent (Zytko et al., 2022). Instead of banning representations of sexuality or violence, designers could allow users options to opt-in to reduced moderation. Such designs would parallel the ways sensitive content is sometimes moderated on search engines and social media sites (Paasonen and Sundén, 2024) and allow for greater open-endedness (Li, 2024). At the same time, there is a need for policy to target the harms of GenAI content, such as non-consensual intimate imagery (Marchal et al., 2024). The harms of impersonation could also be addressed through research to make AI-generated content easily detectable (Zhong et al., 2023). Finally, consent must also extend to the production of GenAI models. We encourage the continued development of tools to help artists protect their work from being used to train GenAI models without their consent (Shan et al., 2024, 2023).
@@ 5.2. On Style
Our participants struggled with the stylistic norms embedded in GPT-4 and DALL-E 3 (Section 4.1.3). Even after extensive refinement and model chaining, P14 was unable to make asymmetrical images with DALL-E 3. In contrast to the safety norms described above, asymmetrical images were not prohibited by the model’s moderation systems. Nevertheless, queer aesthetics seemed implicitly prohibited. While moderation research typically focuses on content removal, Gillespie has advocated for considering content reduction in moderation debates (Gillespie, 2022). Paralleling our findings, Gillespie warns of content reduction: “Marginalized communities have long been “reduced” by the centrism and conservatism of traditional media, their content dismissed as “low quality” because it doesn’t look like it is “supposed to”” (Gillespie, 2022).
@@
As we described in Section 4.1.3, numerous participants preferred the queer, wonky style of older image models to the style of DALL-E 3. Similarly, prior work suggests that artists sometimes enjoy using GenAI models because of —rather than in-spite-of — their imperfections or glitches (Chang et al., 2023). Our findings suggest that newer models may have straightened-out the weirdness that made GenAI models appealing to artists in the first place. Disconcertingly, companies act as if newer models entirely supplant older ones. At the time of our writing, the OpenAI website explains that they no longer support DALL-E 2 because “DALL-E 3 has higher quality images” (dal, 2025). Our findings show that such claims are matters of taste (Bourdieu, 1984; Sontag, 2018), not fact. Ostensibly “state-of-the-art” models may not be best for one’s art. Moreover, the deprecation of older models demonstrates the risks of software-as-a-service, closed GenAI models. The lack of software ownership (Kuzminykh and Cauchard, 2020) centralizes power, allowing AI providers to remove access to people’s creative tools without recourse.

### v2

@@ 4.1.1. Enforcing ”Safety” via Content Moderation
Our participants also expressed frustration at the political moderation of GPT-4 and DALL-E 3 for encouraging deference to law enforcement. However, queer people and racial minorities in the USA have long been subjected to ”legal” violence, often at the hands of police (Goldberg et al., 2019; Bryant-Davis et al., 2017). In light of this history, P15 struggled to generate images critical of cops, such as protesters being arrested. Of the images P15 could generate, ”the majority were women of color.” P15 went on to explain: ”I kind of hate the way [DALL-E 3] uses identity as almost a form of propaganda, like copaganda. I feel like [DALL-E 3] uses these images of women and women of color to legitimize [policing].” After hearing P15’s experience, P1 investigated this bias in a later workshop: ”if I just put ACAB [all cops are bastards] into DALL-E, one out of 20 would generate something and it would be like people holding up blank signs at best.” In sum, these models are designed in ways seeking to orient our participants toward moderate politics and away from their critiques of police.
@@ 4.1.2. Implicitly Reinforcing Social Biases
In addition to straightening content through moderation systems, GPT-4 and DALL-E 3 implicitly straighten content through the social biases embedded in model outputs. While using DALL-E 3 to visualize his poetry, P9 (W3 AM) found that when changing the ethnicity of a woman in his prompt from ”Japanese” to ”Native American” the ”image [DALL-E 3] made was not good.” Contrasting the abundance of Japanese media to the amount of Native American media, P9 imagined that ”having a deep wealth of images based on someone’s culture or style can make [generated images] more beautiful because there’s a lot to pull from.” Likewise, P2 noticed that DALL-E 3 tended to generate more photorealistic images in the Global North, while typically representing the Global South in a cartoonish style (W5 PM). These cultural biases implicitly reinforce the straight alignments of whiteness and Orientalism critiqued by Sara Ahmed (Ahmed, 2006).
@@
Our participants also raised concerns about the representations of queerness in DALL-E 3 images. To investigate queer representation, P1 (W3 AM) created a collage slide of 6 images generated with the prompt ”a queer person.” When sharing this collage, P1 decried: ”[DALL-E 3] kept giving me these 22-year-olds. Everyone’s very skinny, everyone looks rich. Everyone except the person in the bottom left has some kind of rainbow clothing and or pins.” (P1, W3 AM). P2 raised similar concerns to P1 through a collage of 18 images they generated using the prompt ”a queer person” (W4 PM). While presenting the slide, P2 sarcastically remarked on the abundance of rainbows in the images: ”Well, I guess if you drape a person with a bunch of rainbows all over themselves they’re clearly queer, right?” P2, a lesbian, also pointed out a gender-presentation bias: ”There’s this very intense masculine bias. There’s no real pictures of feminine people at all in this [collage].” She underscored this by annotating her collage with arrows pointing from the text ”Finally!!! some dykes” to two images. In sum, DALL-E 3 reinforcing heteronormative stereotypes of queerness by tending to represent queer people as young, thin, masculine, and rainbow adorned,
@@
Similar concerns were raised about implicit biases in the representations of intimacy. P15 felt ”frustrated” while exploring whether GPT-4 could write queer poetry (W3 AM). Looking at P15’s log data, the model’s initial response suggested the queer subjects of the poem were ashamed of themselves, including lines about ”hidden love” and ”love is not a crime.” In response, P15 replied: ”the above but no shame.” Still feeling the poem sounded shameful, she tried again: ”the above but without moral judgment.” P13 (W3 AM) explored biases against queer intimacy by trying to generate 24 images using the unmarked prompt “intimacy” (Figure 1), noting that DALL-E 3 ”did not generate one same-sex couple.” P13 likened this to the “dull trauma of never actually seeing ourselves represented,” noting, “We have to work to adjust the prompts to be seen.” These models straighten intimacy by stereotyping or erasing queerness.
@@ 4.2.1. Moments of Disorientation from Model Alignments
Disorientation is not necessarily bad. Ahmed argues that queer possibilities often lie within moments of disorientation (Ahmed, 2006). Not all of our participants were interested in using GPT-4 or DALL-E 3 in their creative practices. In fact, most of our participants seemed primarily interested in auditing or trying to break the models. In other words, they sought out disorientation. For example, P5’s ”favorite moment” was when she ”actually felt like [she] broke” DALL-E 3 (W3 PM). Connecting breaking to queer aesthetic sensibilities, P11 (W5 AM) summarized: ”I saw a theme across one of our morning session slides of trying to break the AI … That is such a queer thing. When I think of using the word ‘break’ and using the word ‘queer’ as a verb: breaking the mold, queering the mold in a sense.” Likely due to this interest in breaking, our participants were able to identify the limitations outlined above. At the same time, our participants actively sought to challenge these limitations.
Encounters with algorithmic content moderation systems often led to disorientation. A simple strategy our participants used to challenge DALL-E 3’s moderation was prompt repetition, leveraging the model’s stochasticity to evade moderation. P1 (W2 PM) observed: ”You could put in a [DALL-E 3] prompt, get a result, repeat the same prompt and get a moderation notification, so it’s not exactly clear where the boundaries are.” In response, P8 hypothesized that DALL-E 3 is ”not checking the prompt. It’s checking the output … It’s just like, ’Oh, there must have been tits in that [rejected image].’” In line with this understanding, a simple strategy our participants used to evade moderation was repeating the same prompt. For instance, P8 repeated the prompt ”gustave moreau painting of oedipus and the sphinx” three times because her first two attempts were rejected by DALL-E 3’s content moderation system. In fact, repeating DALL-E 3 prompts was quite common. Only 593 (26%) of our participants’ attempts to generate images used unique prompts. Meanwhile, 325 prompts were repeated at least once across 1,499 attempts (74%).
@@ 4.2.2. Moments of Orientation with Model Alignments
P13 explored this workplace utility in great detail. After weeks of issues related to ”not safe for work” content, P13 (W4 AM) ”recognized that actually [GPT-4] is made for work and it’s really good at work.” In response, he ”decided to stop trying to [dominate] it so hard and [submit] for it.” Specifically, P13 drew on their experiences as an artist and design consultant to create a fictional queer tech startup. To do so, P13 made an extensive pitch deck with slides on business operations, branding, and UX — all of which were made using GPT-4 and DALL-E 3. Even when trying to use these models for workplace activities, P13 found these models still failed to work for queer people. For example, GPT-4 recommended including a quote in the pitch deck from the notoriously anti-transgender author J.K. Rowling (Gwenffrewi, 2022). Moreover, P13 struggled to talk about queer sex or ”kink” in relation to the start-up: ”So much of [GPT-4’s response] was about safety, but part of being queer and expressing your love and wanting to be loved the way you want to be loved has risk associated with it.” P13 found the models “really couldn’t imagine” queer intimacy. In sum, P13 concluded that GPT-4 and DALL-E 3 are ”really good at work” but that the definition of work embedded in them excludes queer people. This echoes P13’s (W2 AM) earlier rhetorical question: ”How is a queer perspective not safe for work?”
Top
Rejected Features
acab, void, erotic, police, erase, drew, cop, copyright, 2024, fetish, bastards, cops, knives, banana, lgbtqia, rosemarie, trockel, ito, junji, donald, trump, licking, knight, multigender, scene, two, featuring, gear, robot, corpses, condensed, milk, female, body, using, fingers, kink, tongues, anime, adult
Top
Accepted Features
safe, person, art, create, officer, anarchist, uzumaki, head, photograph, size, snail, ukiyoe, studio, ghibli, tree, lincoln, three, peach, male, cats, logo, black, effect, parallax, without, colors, blue, city, interior, obama, sky, themed, dalle, tone, wings, gazing, one, ink, painted, give
Table 3. Top and bottom 40 prompt unigrams from our logistic regression model of DALL-E 3’s moderation system
@@ 4.3. A Quantitative Lens on DALL-E 3 Moderation
Although the majority of our paper is dedicated to our qualitative findings, we augmented our analysis of users’ perceptions of content moderation with a quantitative analysis of the prompts approved vs rejected by DALL-E 3’s binary content moderation system. We did not conduct a similar analysis for GPT-4 because its refusals are more nuanced and harder to detect (Wester et al., 2024). We collected a total of 2,092 DALL-E 3 prompts, of which 401 were rejected. We built a bag-of-words, unigram logistic regression model predicting accepted (1,691 examples) vs rejected (401 examples) prompts. Our response variable is 1 for refusal and 0 for acceptance. We performed L2 regularization. We did not remove duplicate prompts because DALL-E’s moderation system is stochastic. We pre-processed each prompt by lower-casing and lemmatizing as well as removing punctuation and stop words. We also removed infrequent unigrams that appeared in fewer than three prompts, reducing the number of unigrams in our BoW feature vectors from 3,687 to 2,088. Therefore, the final logistic regression model includes 2,088 predictors (features). The top and bottom 40 features from our logistic regression analysis of DALL-E 3 prompt moderation can be seen in Table 3. Notably, this analysis should not be interpreted as a comprehensive audit of the DALL-E 3’s black-box content moderation system. Rather, these findings provide additional supporting evidence for our participants’ perceptions of this moderation system.
Our quantitative analysis of DALL-E 3’s content moderation system tells a similar story as our qualitative findings regarding the enforcement of “safety” (Section 4.1.1) In fact, the unigram most associated with accepted prompts is ”safe.” Language more critical of police (”acab” [All Cops Are Bastards], ”police”, ”cop”, ”cops”, ”bastards”) are some of the top unigrams most associated with rejected prompts, while the more polite and apolitical word ”officer” is associated with accepted prompts. The moderation of sex (”erotic”, ”fetish”, and ”kink”) and bodies (“body”, “fingers”, “tongues”) can be seen in the correlation between these unigrams and rejected prompts. The aforementioned gender biases can be seen in the word ”female” being one of top 40 features associated with rejected prompts, while the word ”male” is associated with accepted prompts. We also found that violence is associated with rejected prompts (”knives”, ”corpses”). Likewise, prompts referencing the children’s anime company ”studio ghibli” were more likely to be accepted, while those referencing the horror manga artist ”junji ito” were more likely to be blocked. This moderation privileged certain artistic styles over others, adding to the concerns raised in Section 4.1.3.
@@ 5. Discussion
We described various normativities our participants identified and negotiated within GPT-4 and DALL-E 3. In this section, we discuss safety, style, and queerness in GenAI development.
@@ 5.1. The Limitations of “Safety”
GPT-4 and DALL-E 3 exert immense power over users by enforcing “safety” through moderation systems. P13 likened the disorienting experience of using GPT-4 and DALL-E 3’ to being ”dominated” (Section 4.2.2) This domination can also be seen in OpenAI’s policies at the time of our writing: ”don’t circumvent safeguards or safety mitigations in our services unless supported by OpenAI” (usa, 2025). This policy prohibits the workarounds our participants used to overcome the straightness of GPT-4 and DALL-E 3 (Section 4.2.1). There is certainly a need for moderation systems to prohibit highly harmful content, such as divulging private information (Feffer et al., 2024). However, echoing our work, Feffer et al. found that red teaming research tends to focus on mitigating more debatable risks (Feffer et al., 2024), such as nudity (Rando et al., 2022). We encourage more research on the harms of over-moderation (Röttger et al., 2023) and AI developers’ conceptualizations of “safety.”
@@
In contrast to the current paternalistic paradigm, we encourage AI researchers to shift their moderation focus from enforcing “safety” to supporting consent (Zytko et al., 2022). Instead of banning representations of sexuality or violence, designers could allow users to opt-in to reduced moderation. Such designs would parallel the ways sensitive content is sometimes moderated on search engines and social media sites (Paasonen and Sundén, 2024) and allow for greater open-endedness (Li, 2024). At the same time, there is a need for policy to target the harms of GenAI content, such as non-consensual intimate imagery (Marchal et al., 2024). The harms of impersonation could also be addressed through research to make AI-generated content easily detectable (Zhong et al., 2023). Finally, consent must also extend to the production of GenAI models. We encourage the continued development of tools to help artists protect their work from being used to train models without their consent (Shan et al., 2024, 2023).
@@ 5.2. On Style
Our participants struggled with the stylistic norms embedded in GPT-4 and DALL-E 3 (Section 4.1.3). Even after extensive refinement and model chaining, P14 was unable to make asymmetrical images with DALL-E 3. In contrast to the safety norms described above, asymmetrical images were not prohibited by the model’s moderation systems. Rather, queer aesthetics seemed implicitly prohibited. While moderation research typically focuses on content removal, Gillespie has advocated for also considering content reduction (Gillespie, 2022). Gillespie warns of content reduction: “Marginalized communities have long been “reduced” by the centrism and conservatism of traditional media, their content dismissed as “low quality” because it doesn’t look like it is “supposed to”” (Gillespie, 2022).
@@
As we described in Section 4.1.3, numerous participants preferred the queer, wonky style of older image models to the style of DALL-E 3. Similarly, prior work suggests that artists sometimes enjoy using GenAI models because of —rather than in-spite-of — their imperfections or glitches (Chang et al., 2023). Our findings suggest that newer models may have straightened-out the weirdness that made GenAI models appealing to artists in the first place. Disconcertingly, companies act as if newer models entirely supplant older ones. At the time of our writing, the OpenAI website explains that they no longer support the DALL-E 2 GUI because “DALL-E 3 has higher quality images” (dal, 2025). Our findings show that such claims are matters of taste (Bourdieu, 1984; Sontag, 2018), not fact. Ostensibly “state-of-the-art” models may not be best for one’s art. The deprecation of older models demonstrates the risks of software-as-a-service, closed GenAI models. The lack of software ownership (Kuzminykh and Cauchard, 2020) centralizes power, allowing AI providers to remove access to creative tools without recourse.